OpenStudy (anonymous):

find the value of 12÷4×3+5−2

OpenStudy (anonymous):

12/4 * 3 + 5 - 2 = 3 * 3 + 5 - 2 = 9 + 5 - 2 = 14 - 2 = 12

OpenStudy (anonymous):

Your previous question was not very clear.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

12